OpenStudy (anonymous):

The following is a list of the number of siblings that various students have: 1, 1, 3, 0, 4, 1, 2, 1, 2, 0, 3, 2, 1, 5, 4, 2. What is the mean?

OpenStudy (anonymous):

1 2 4 5 answers

OpenStudy (campbell_st):

well the sum of the scores is 32 and the number of scores is 16 so \[Mean = \frac{Sum..of..scores}{number..of..scores}\]

OpenStudy (anonymous):

Mean= sum of elements in the sets/ number of elements in the set Mean= 1+1+ 3+ 0+ 4+1+ 2+ 1+2+0+ 3+ 2+1+ 5+ 4+ 2/16 Mean=32/16 Mean=2 So, option number B is the answer :)